OpenStudy (anonymous):

whats 14*8/2-9*square root of 9

OpenStudy (jackellyn):

9*sqrt 9 = 9(3) = 27 14*8/2 =56 56-27=29

OpenStudy (anonymous):

yu sure?

OpenStudy (jackellyn):

should be right

OpenStudy (anonymous):

okay thanks ;)